300-volt power station is connected to a factory through two power lines. If the factory is 5 Km away from the station and electric potential difference at factory 220 V, current intensity 100 A, calculate: a) The resistance of unit length. b) Radius of CSA of the wire. given that its resistivity 1.5x10-8 2.m.
1
Expert's answer
2021-09-03T12:12:25-0400

The potential drop on the power line is equal

"\Delta V=300\:\rm V-220\: V=80\: V"

Hence, the resistance of the power line

"R=\frac{\Delta V}{I}=\frac{80\:\rm V}{100\:\rm A}=0.8\:\Omega"

a) The resistance of unit length

"R/l=0.8\:\Omega/(2*5000\:\rm m)=8*10^{-5}\:\Omega/m"

b) Radius of CSA of the wire

"r=\sqrt{\frac{A}{\pi}}=\sqrt{\frac{\rho l/R}{\pi}}\\
=\sqrt{\frac{1.5*10^{-8}\:\Omega\cdot\rm m}{3.14*8*10^{-5}\:\rm\Omega/m}}=7.7*10^{-3}\:\rm m=7.7\: mm"